# DiffDecompose: Layer-Wise Decomposition of Alpha-Composited Images via Diffusion Transformers

### v2

@@ 1 Introduction
With the development of diffusion models, image-to-image [42, 16, 62, 65] and text-to-image [6, 18, 10] generation methods have been extensively studied and have a wide applications in image inpainting [12, 49, 61, 59], video generation [5, 58, 76, 60], and image segmentation [36, 2, 22] to obtain high-quality images and user-friendly interfaces. Most existing approaches rely on mask-based constraints to guide the model toward modifying or preserving specific image regions, or assume temporal consistency by treating moving object information as static to infer future or missing frames. Despite their success, these methods face two major limitations when applied to more complex and realistic scenarios. Firstly, the assumption of known separation regions or static objects is too simple, making it difficult to apply to more complex scenarios. Secondly, the noise in the diffusion model’s forward phase may destroy the image’s original information, reducing the image context’s usability.
@@
Recently, a new type of method, termed layer-wise decomposition [53, 64, 71, 30, 13, 3, 21], has gradually entered the field of computer vision and become a potential method to solve the problem of these generative models. The core objective of layer-wise decomposition is to decompose a single composited image into its constituent layers, each containing the foreground object, its associated alpha matte (transparency), and potential depth ordering. Such a decomposition enables more granular control over individual layers, allowing for other tasks such as image inpainting [44], matting [35], scene understanding [66], and content creation [52]. Nonetheless, these decomposing methods will fail to recover the accurate background when the foreground itself is layer-level semi-transparent or transparent, since they still rely on region-based assumptions. Most importantly, current methods, such as image inpainting [69, 39, 17, 26] and layer decomposition [71, 64], cannot generalize to the real world to tackle arbitrary alpha-blended images with diverse contents and complex layer structures.
@@
In this paper, our goal is to delve into a novel task that focuses on semi-transparent or transparent layer-wise decomposition. As shown in Figure 1 (a), this novel task does not require any mask-based information for background assumption and no clear visual contrast to provide depth information, and under such conditions, a natural question is how to decompose these transparent or semi-transparent layers from the background ? Naturally, a single composited image often corresponds to multiple plausible decompositions due to the complex entanglement of color and transparency across layers, leading to additional difficulties in reconstructing the original scene. Therefore, this proposed task faced three main challenges not addressed in prior works: (1) Layer ambiguity and coupling of color and transparency: foreground and background occupy the same visual plane and lack the separability via depth or edge contrast, thereby introducing significant layer ambiguity. Furthermore, critical foreground details may be partially embedded in the alpha channel, rendering them inaccessible through RGB information alone. (2) Generalization in real-world scenarios: The above methods directly remove the pixel values of a specific area and use the network to generate new pixel values based on background reasoning. Such a manner largely overlooks the exploration of pixel-level correlations between the occluded area and the background, leading to the failure of generalization in various realistic scenarios. (3) Lack of large-scale dataset: A significant challenge in layer decomposition is the lack of large-scale and high-quality datasets. Current datasets [53, 64] are often generated from AI, and the final overlapped images tend to have a pixel-level difference from the initial foreground and background, resulting in the failure of layer-wise decomposition.
@@ 2 Related Work
Diffusion Model. Diffusion Models (DMs) [46, 43] have significantly advanced image generation and editing, surpassing GANs [14, 34, 38, 34, 33, 55] in tasks such as artistic synthesis [9], instruction-guided transformation [7], and controllable editing [25, 68]. Recent extensions have explored applications in 3D modeling [31] and scene-consistent generation [37], showcasing the broad generative capacity of diffusion models. Despite these advances, diffusion models remain an open and ill-posed challenge, particularly in scenarios involving transparency, semi-transparency, and nonlinear blending. In this work, we present a novel generative task, namely Layer-Wise Decomposition of Alpha-Composited Images, and propose a diffusion-based framework DiffDecompose, to decompose complex semi-transparent/transparent layer-level composited images into plausible foreground and background layers. To support this novel task, we construct a first, large-scale and high-quality dataset AlphaBlend, which mainly contains layer-level semi-transparent or transparent images.
@@
Controllable Generation. Recent advances in controllable diffusion models [72, 67, 39, 40, 60] have enabled powerful image editing capabilities, particularly in inpainting. Existing approaches are typically categorized into text-based and mask-based methods. Text-based methods [70, 77, 61] rely on prompt engineering to achieve multi-task unification, but their performance is highly sensitive to prompt accuracy, often resulting in degraded outputs when instructions are imprecise [63]. Mask-based methods [69, 39, 17] focus on preserving background integrity via accurate region masks, yet they introduce noise to masked areas and perform iterative denoising, which weakens the conditioning signal and limits performance in complex blending scenarios. To this end, we extend a noise-free conditioning paradigm inspired by OminiControl [50], from a novel point to reformulate the object occlusion task [70, 17, 49] or layer decomposition task [53, 64] as a probabilistic generation task, where the goal is to model the posterior distribution over plausible layer decompositions conditioned on the observed composite image. Under this definition, not only can we solve the common object-level or channel-level occlusion task, but also extend this idea to our proposed layer-level semi-transparent/transparent task.
@@
In-context Learning. In-context learning, originally developed in large language models [8, 74], has recently been extended to vision tasks such as object editing and region manipulation [74, 54, 47, 4, 73, 11, 48, 20]. While prior works leverage contextual cues for object insertion or style transfer, none have addressed layer-wise decomposition, particularly in semi-transparent or transparent settings. In this paper, we introduce In-Context Decomposition (ICD), which enables the model to infer multiple compositional layers based on contextual prompts without explicit supervision. Furthermore, we propose a Layer Position Encoding Cloning (LPEC) strategy to preserve spatial alignment and reinforce information consistency between the decomposed layers and the original image.
@@ 3.2 Problem Formulation and Overview
Problem Formulation. As shown in blue background of Figure 2, to obtain the background from the observed image $\mathbf{z}\in\mathbb{R}^{H\times W\times 3}$, the current method [69, 39, 17] regards this problem as an image edition task or a common mask-based layer-wise decomposition task, which can be written as:
[…]
@@
[…]
@@
Framework Overview. As shown in Figure 3, the overview architecture of DiffDecompose can be divided into two steps: (1) Utilize the frozen VAE [27] to respectively encode a semi-transparent/transparent foreground, a common RGB background, and a condition alpha-blend image, to obtain their latent spatial-level features. In the meantime, the prompts are fed into the frozen T5XXL [41] to extract the text-level features. (2) The process of decomposition is realized by In-context Decomposition (ICD) by concatenating the clean condition token and noise tokens, and utilizing the bidirectional attention to achieve conditional generation. During the process of ICD, we also devise a Layer Position Encoding Clone strategy to achieve consistency between the specific layer and the original layer information, and utilize the MMA for the interaction between different modalities.
@@ 3.3 In-Context Decomposition
This operation ensures that both the background and the composite tokens share a unified coordinate frame, i.e., $\forall(i,j),\quad\tilde{c}_{y}^{(i,j)}-\tilde{c}_{z}^{(i,j)}=c_{y}^{(i,j)}-c_{z}^{(i,j)}$ preserves relative positional structure during attention computation. In contrast, the foreground tokens $c_{x}$ are excluded from this positional cloning, i.e.,$\tilde{c}_{x}=c_{x}$. This separation avoids spatial entanglement between layers and prevents unwanted feature blending. In transparent or semi-transparent compositions, foreground and background often overlap spatially but differ in semantic meaning and transparency. Maintaining disjoint positional encodings: $\mathrm{PE}_{x}\perp\mathrm{PE}_{z}$, the model prevents spatial interference between foreground and background layers. By maintaining a distinct positional space for the foreground, the model is encouraged to preserve its semantic and spatial independence, thereby enhancing the fidelity of layer disentanglement in semi-transparent or fully transparent scenes.
Multi-Modality Attention. After applying LPEC, the latent tokens $c_{x}$, $c_{y}$, and $c_{z}$ are then concatenated along the sequence dimension to perform joint attention. The Multi-modal attention mechanisms [18] are utilized to provide conditional information for the denoising of the alpha-composed image. Furthermore, by maintaining $y$ in a noise-free state, we ensure the retention of high-frequency textures and fine structural details from the original image, thereby preventing degradation during iterative denoising. This process can be written as:
@@
where $Q,K,V$ are the query, key, and value projections of the concatenated sequence $[\tilde{c}_{z};\tilde{c}_{x};\tilde{c}_{y};c_{T}]$, $d$ is the feature dimension, and $c_{T}$ represents the textual condition tokens extracted from a task-specific prompt. In particular, the description consists of the foreground image $x$, the background image $y$, and their composited observation $z$, (e.g., “Three sub-images. <image-1>: a transparent glass, <image-2>: this is a dinning room with some chairs, <image-3>: The overlapped of <image-1> and <image-2>.), which guide the semantic understanding of the inter-layer relationships.
@@ 4.1 Experimental Settings
Implementation Details. We initialize our model using the pre-trained parameters of the Flux.1 [28] architecture and fine-tune it across six representative subtasks. The model is trained using a LoRA [18] rank of 128, batch size of 1, and a learning rate of ${10}^{-4}$ on a single H20 GPU for 30,000 steps. For the subtasks of translucent flare removal, translucent occlusion removal, semi-transparent watermark and removal, all input images are resized to 512×512. For the subtasks of transparent glassware decomposition, semi-transparent cell decomposition, and X-ray contraband decomposition, input images are resized to 512×1024, and the training is conducted for 50,000 steps under the same hyperparameter configuration unless otherwise specified.
Baseline. We compare our method with the state-of-the-art inpainting methods, which include the Inpainting Anything [69], diffusion-based models SDXL-Inpaint [39], Flux-control-Inpaint [51], PowerPaint [77] and ClipAway [17]. To ensure fair comparison and reduce inconsistencies or boundary artifacts, we apply mask dilation for five iterations using a 5×5 kernel prior to inpainting.
Benchmark. To provide a rigorous and comprehensive evaluation, we evaluate our DiffDecompose and baselines on two datasets: our proposed AlphaBlend dataset and the public LOGO dataset [15]. For our proposed AlphaBlend dataset (described in Sec. 3.1), DiffDecompose infers its six scenarios, including: X-ray contraband decomposition, transparent glassware decomposition, semi-transparent cell decomposition, semi-transparent watermark removal, translucent occlusion removal, and translucent flare removal, with 300-500 test images. For LOGO evaluation, we utilize its three test datasets, LOGO-H, LOGO-L, and LOGO-G.
@@ 4.2 Main Results
PowerPoint [77] 10.0330 0.9726 0.0360 34.64 6.4480 0.9870 0.0212 30.49 8.1397 0.9811 0.0272 32.42
Flux-ControlNet [51] 11.7016 0.9640 0.0508 64.67 8.7095 0.9800 0.0351 61.22 14.5943 0.9670 0.0705 37.28
SDXL-inpainting [39] 8.1936 0.9781 0.0302 35.24 4.9184 0.9899 0.0172 30.57 6.3711 0.9848 0.0221 32.37
ClipAaway [17] 11.3593 0.9690 0.0463 36.55 7.5353 0.9851 0.0275 32.40 9.3021 0.9785 0.0356 33.21
Inpaint Anything [69] 4.8252 0.9849 0.0143 27.50 3.0461 0.9931 0.0072 23.71 3.8823 0.9898 0.0101 25.90
@@
PowerPoint [77] 31.4009 0.8808 0.0571 59.70 10.6886 0.9727 0.0236 33.35 9.55 0.8504 0.1526 36.46
Flux-ControlNet [51] 35.9988 0.8430 0.1848 82.21 17.3372 0.9551 0.0451 68.41 12.08 0.9701 0.0407 89.47
SDXL-inpainting [39] 21.7313 0.9163 0.0923 50.89 14.4375 0.9596 0.0421 73.36 9.70 0.9764 0.0274 79.18
ClipAaway [17] 35.3515 0.8641 0.1561 60.80 13.9394 0.9649 0.0333 47.64 7.58 0.9590 0.0180 49.79
Inpaint Anything [69] 11.5273 0.9527 0.0484 34.58 8.2335 0.9775 0.0189 31.30 3.3028 0.9862 0.0093 36.42
@@
PowerPoint [77] 7.1830 0.9826 0.0134 39.90 90.4829 0.1319 0.8150 404.70 89.1031 0.1336 0.8092 421.63
Flux-ControlNet [51] 7.6561 0.9808 0.1550 45.49 96.0870 0.2158 0.8016 253.40 87.8764 0.1991 0.8126 239.67
SDXL-inpainting [39] 7.0060 0.9845 0.0125 39.38 63.1518 0.3539 0.7269 229.96 64.6848 0.3296 0.7351 209.20
ClipAaway [17] 9.9123 0.9789 0.0200 55.92 97.5711 0.2416 0.8362 255.26 96.2468 0.2344 0.8427 235.00
Inpaint Anything [69] 4.8260 0.9856 0.0080 38.13 83.3769 0.3058 0.9040 462.58 83.2172 0.3090 0.9046 477.38
@@
In complex real-world scenarios such as X-ray contraband removal and translucent window occlusions, our model achieves up to 4× lower RMSE and 3× better LPIPS than inpainting-based baselines, which struggle with color entanglement and global transparency. In particular, for challenging global occlusions like "Rain" and "Light", our method significantly surpasses all baselines, while others fail without explicit masks. Although our FID in LOGO-L is slightly higher than Inpaint Anything [69] (24.20 vs 23.71), which is attributed to the highly structured and mask-friendly layout of LOGO-L, providing strong mask priors. Importantly, our model is mask-free and shows consistently better structural accuracy across all LOGO datasets (e.g., RMSE ↓1.6, SSIM ↑0.002, LPIPS ↓0.011 over Inpaint Anything on LOGO-H). These results confirm that DiffDecompose not only generalizes across our six diverse datasets but also outperforms existing methods on public benchmarks, without requiring additional spatial priors.

### v3

@@ 1 Introduction
With the development of diffusion models, image-to-image richardson2021encoding; ding2024enhance; xu2024ufogen; yang2025unified and text-to-image betker2023improving; esser2024scaling; chen2023pixart generation methods have been extensively studied and have a wide applications in image inpainting chen2024zero; sun2023alphaclip; xie2022smartbrush; wu2024towards, video generation bar2024lumiere; wu2023tune; zhou2022magicvideo; wu2024motionbooth, and image segmentation minaee2021image; amit2021segdiff; gu2024diffusioninst to obtain high-quality images and user-friendly interfaces. Most existing approaches rely on mask-based constraints to guide the model toward modifying or preserving specific image regions, or assume temporal consistency by treating moving object information as static to infer future or missing frames. Despite their success, these methods face two major limitations when applied to more complex and realistic scenarios. Firstly, the assumption of known separation regions or static objects is too simple, making it difficult to apply to more complex scenarios. Secondly, the noise in the diffusion model’s forward phase may destroy the image’s original information, reducing the image context’s usability.
@@
Recently, a new type of method, termed layer-wise decomposition tudosiu2024mulan; yang2024generative; zhang2024transparent; lee2024generative; chen2025transanimate; bai2025layer; gu2023factormatte, has gradually entered the field of computer vision and become a potential method to solve the problem of these generative models. The core objective of layer-wise decomposition is to decompose a single composited image into its constituent layers, each containing the foreground object, its associated alpha matte (transparency), and potential depth ordering. Such a decomposition enables more granular control over individual layers, allowing for other tasks such as image inpainting shi2024foodfusion, matting lu2025efficient, scene understanding yang2024layerpano3d, and content creation xie2025anywhere. Nonetheless, these decomposing methods will fail to recover the accurate background when the foreground itself is layer-level semi-transparent or transparent, since they still rely on region-based assumptions. Most importantly, current methods, such as image inpainting yu2023inpaint; sdxl; ekin2024clip; jiang2025smarteraserremoveimagesusing and layer decomposition zhang2024transparent; yang2024generative, cannot generalize to the real world to tackle arbitrary alpha-blended images with diverse contents and complex layer structures.
@@
In this paper, our goal is to delve into a novel task that focuses on semi-transparent or transparent layer-wise decomposition. As shown in Figure 1 (a), this novel task does not require any mask-based information for background assumption and no clear visual contrast to provide depth information, and under such conditions, a natural question is how to decompose these transparent or semi-transparent layers from the background ? Naturally, a single composited image often corresponds to multiple plausible decompositions due to the complex entanglement of color and transparency across layers, leading to additional difficulties in reconstructing the original scene. Therefore, this proposed task faced three main challenges not addressed in prior works: (1) Layer ambiguity and coupling of color and transparency: foreground and background occupy the same visual plane and lack the separability via depth or edge contrast, thereby introducing significant layer ambiguity. Furthermore, critical foreground details may be partially embedded in the alpha channel, rendering them inaccessible through RGB information alone. (2) Generalization in real-world scenarios: The above methods directly remove the pixel values of a specific area and use the network to generate new pixel values based on background reasoning. Such a manner largely overlooks the exploration of pixel-level correlations between the occluded area and the background, leading to the failure of generalization in various realistic scenarios. (3) Lack of large-scale dataset: A significant challenge in layer decomposition is the lack of large-scale and high-quality datasets. Current datasets tudosiu2024mulan; yang2024generative are often generated from AI, and the final overlapped images tend to have a pixel-level difference from the initial foreground and background, resulting in the failure of layer-wise decomposition.
@@ 2 Related Work
Diffusion Model. Diffusion Models (DMs) song2020denoising; rombach2021highresolution have significantly advanced image generation and editing, surpassing GANs chen2019improved; liu2018image; pathak2016context; liu2018image; liu2022partial; wang2024semflow in tasks such as artistic synthesis chen2024inverse, instruction-guided transformation brooks2023instructpix2pix, and controllable editing hui2024; yu2025. Recent extensions have explored applications in 3D modeling li2025garment and scene-consistent generation pan2025model, showcasing the broad generative capacity of diffusion models. Despite these advances, diffusion models remain an open and ill-posed challenge, particularly in scenarios involving transparency, semi-transparency, and nonlinear blending. In this work, we present a novel generative task, namely Layer-Wise Decomposition of Alpha-Composited Images, and propose a diffusion-based framework DiffDecompose, to decompose complex semi-transparent/transparent layer-level composited images into plausible foreground and background layers. To support this novel task, we construct a first, large-scale and high-quality dataset AlphaBlend, which mainly contains layer-level semi-transparent or transparent images.
@@
Controllable Generation. Recent advances in controllable diffusion models zhang2023adding; ye2023ip-adapter; sdxl; peebles2023; wu2024motionbooth have enabled powerful image editing capabilities, particularly in inpainting. Existing approaches are typically categorized into text-based and mask-based methods. Text-based methods yu2024promptfix; zhuang2024task; xie2022smartbrush rely on prompt engineering to achieve multi-task unification, but their performance is highly sensitive to prompt accuracy, often resulting in degraded outputs when instructions are imprecise xu2025pixel. Mask-based methods yu2023inpaint; sdxl; ekin2024clip focus on preserving background integrity via accurate region masks, yet they introduce noise to masked areas and perform iterative denoising, which weakens the conditioning signal and limits performance in complex blending scenarios. To this end, we extend a noise-free conditioning paradigm inspired by OminiControl tan2025, from a novel point to reformulate the object occlusion task yu2024promptfix; ekin2024clip; sun2023alphaclip or layer decomposition task tudosiu2024mulan; yang2024generative as a probabilistic generation task, where the goal is to model the posterior distribution over plausible layer decompositions conditioned on the observed composite image. Under this definition, not only can we solve the common object-level or channel-level occlusion task, but also extend this idea to our proposed layer-level semi-transparent/transparent task.
@@
In-context Learning. In-context learning, originally developed in large language models mann2020language; zhang2025incontextedit, has recently been extended to vision tasks such as object editing and region manipulation zhang2025incontextedit; wang2025explore; song2025insert; bai2024sequential; zhang2023makes; chen2025edittransfer; song2025makeanything; fang2023explore. While prior works leverage contextual cues for object insertion or style transfer, none have addressed layer-wise decomposition, particularly in semi-transparent or transparent settings. In this paper, we introduce In-Context Decomposition (ICD), which enables the model to infer multiple compositional layers based on contextual prompts without explicit supervision. Furthermore, we propose a Layer Position Encoding Cloning (LPEC) strategy to preserve spatial alignment and reinforce information consistency between the decomposed layers and the original image.
@@ 3.2 Problem Formulation and Overview
Problem Formulation. As shown in blue background of Figure 2, to obtain the background from the observed image $\mathbf{z}\in\mathbb{R}^{H\times W\times 3}$, the current method yu2023inpaint; sdxl; ekin2024clip regards this problem as an image edition task or a common mask-based layer-wise decomposition task, which can be written as:
$p_{\theta}(\mathbf{y}\mid\mathbf{z},\mathbf{m})=\int p_{\theta}(\mathbf{y}\mid\mathbf{z},\mathbf{m})p_{\theta}\left(\mathbf{z}_{0}\mid\mathbf{z_{t}},\mathbf{m}\right)d\mathbf{z}_{0},$ (1)
@@
$p_{\theta}(\mathbf{x},\mathbf{y}\mid\mathbf{z},\tau)=\int p_{\theta}(\mathbf{x},\mathbf{y}\mid\mathbf{z},\tau)p_{\theta}\left(\mathbf{z}_{0}\mid\mathbf{z},\tau\right)d\mathbf{z}_{0}.$ (2)
@@
Framework Overview. As shown in Figure 3, the overview architecture of DiffDecompose can be divided into two steps: (1) Utilize the frozen VAE kingma2013auto to respectively encode a semi-transparent/transparent foreground, a common RGB background, and a condition alpha-blend image, to obtain their latent spatial-level features. In the meantime, the prompts are fed into the frozen T5XXL raffel2020exploring to extract the text-level features. (2) The process of decomposition is realized by In-context Decomposition (ICD) by concatenating the clean condition token and noise tokens, and utilizing the bidirectional attention to achieve conditional generation. During the process of ICD, we also devise a Layer Position Encoding Clone strategy to achieve consistency between the specific layer and the original layer information, and utilize the MMA for the interaction between different modalities.
@@ 3.3 In-Context Decomposition
[…]
Multi-Modality Attention. After applying LPEC, the latent tokens $c_{x}$, $c_{y}$, and $c_{z}$ are then concatenated along the sequence dimension to perform joint attention. The Multi-modal attention mechanisms esser2024scaling are utilized to provide conditional information for the denoising of the alpha-composed image. Furthermore, by maintaining $y$ in a noise-free state, we ensure the retention of high-frequency textures and fine structural details from the original image, thereby preventing degradation during iterative denoising. This process can be written as:
@@
where $Q,K,V$ are the query, key, and value projections of the concatenated sequence $\tilde{c}_{z};\tilde{c}_{x};\tilde{c}_{y};c_{T}]$, $d$ is the feature dimension, and $c_{T}$ represents the textual condition tokens extracted from a task-specific prompt. In particular, the description consists of the foreground image $x$, the background image $y$, and their composited observation $z$, (e.g., “Three sub-images. <image-1>: a transparent glass, <image-2>: this is a dinning room with some chairs, <image-3>: The overlapped of <image-1> and <image-2>.), which guide the semantic understanding of the inter-layer relationships.
@@ 4.1 Experimental Settings
Implementation Details. We initialize our model using the pre-trained parameters of the Flux.1 flux2024 architecture and fine-tune it across six representative subtasks. The model is trained using a LoRA esser2024scaling rank of 128, batch size of 1, and a learning rate of ${10}^{-4}$ on a single H20 GPU for 30,000 steps. For the subtasks of translucent flare removal, translucent occlusion removal, semi-transparent watermark and removal, all input images are resized to 512×512. For the subtasks of transparent glassware decomposition, semi-transparent cell decomposition, and X-ray contraband decomposition, input images are resized to 512×1024, and the training is conducted for 50,000 steps under the same hyperparameter configuration unless otherwise specified.
Baseline. We compare our method with the state-of-the-art inpainting methods, which include the Inpainting Anything yu2023inpaint, diffusion-based models SDXL-Inpaint sdxl, Flux-control-Inpaint Controlnet, PowerPaint zhuang2024task and ClipAway ekin2024clip. To ensure fair comparison and reduce inconsistencies or boundary artifacts, we apply mask dilation for five iterations using a 5×5 kernel prior to inpainting.
Benchmark. To provide a rigorous and comprehensive evaluation, we evaluate our DiffDecompose and baselines on two datasets: our proposed AlphaBlend dataset and the public LOGO dataset cun2021split. For our proposed AlphaBlend dataset (described in Sec. 3.1), DiffDecompose infers its six scenarios, including: X-ray contraband decomposition, transparent glassware decomposition, semi-transparent cell decomposition, semi-transparent watermark removal, translucent occlusion removal, and translucent flare removal, with 300-500 test images. For LOGO evaluation, we utilize its three test datasets, LOGO-H, LOGO-L, and LOGO-G.
@@ 4.2 Main Results
PowerPoint zhuang2024task 10.0330 0.9726 0.0360 34.64 6.4480 0.9870 0.0212 30.49 8.1397 0.9811 0.0272 32.42
Flux-ControlNet Controlnet 11.7016 0.9640 0.0508 64.67 8.7095 0.9800 0.0351 61.22 14.5943 0.9670 0.0705 37.28
SDXL-inpainting sdxl 8.1936 0.9781 0.0302 35.24 4.9184 0.9899 0.0172 30.57 6.3711 0.9848 0.0221 32.37
ClipAaway ekin2024clip 11.3593 0.9690 0.0463 36.55 7.5353 0.9851 0.0275 32.40 9.3021 0.9785 0.0356 33.21
Inpaint Anything yu2023inpaint 4.8252 0.9849 0.0143 27.50 3.0461 0.9931 0.0072 23.71 3.8823 0.9898 0.0101 25.90
@@
PowerPoint zhuang2024task 31.4009 0.8808 0.0571 59.70 10.6886 0.9727 0.0236 33.35 9.55 0.8504 0.1526 36.46
Flux-ControlNet Controlnet 35.9988 0.8430 0.1848 82.21 17.3372 0.9551 0.0451 68.41 12.08 0.9701 0.0407 89.47
SDXL-inpainting sdxl 21.7313 0.9163 0.0923 50.89 14.4375 0.9596 0.0421 73.36 9.70 0.9764 0.0274 79.18
ClipAaway ekin2024clip 35.3515 0.8641 0.1561 60.80 13.9394 0.9649 0.0333 47.64 7.58 0.9590 0.0180 49.79
Inpaint Anything yu2023inpaint 11.5273 0.9527 0.0484 34.58 8.2335 0.9775 0.0189 31.30 3.3028 0.9862 0.0093 36.42
@@
PowerPoint zhuang2024task 7.1830 0.9826 0.0134 39.90 90.4829 0.1319 0.8150 404.70 89.1031 0.1336 0.8092 421.63
Flux-ControlNet Controlnet 7.6561 0.9808 0.1550 45.49 96.0870 0.2158 0.8016 253.40 87.8764 0.1991 0.8126 239.67
SDXL-inpainting sdxl 7.0060 0.9845 0.0125 39.38 63.1518 0.3539 0.7269 229.96 64.6848 0.3296 0.7351 209.20
ClipAaway ekin2024clip 9.9123 0.9789 0.0200 55.92 97.5711 0.2416 0.8362 255.26 96.2468 0.2344 0.8427 235.00
Inpaint Anything yu2023inpaint 4.8260 0.9856 0.0080 38.13 83.3769 0.3058 0.9040 462.58 83.2172 0.3090 0.9046 477.38
@@
In complex real-world scenarios such as X-ray contraband removal and translucent window occlusions, our model achieves up to 4× lower RMSE and 3× better LPIPS than inpainting-based baselines, which struggle with color entanglement and global transparency. In particular, for challenging global occlusions like "Rain" and "Light", our method significantly surpasses all baselines, while others fail without explicit masks. Although our FID in LOGO-L is slightly higher than Inpaint Anything yu2023inpaint (24.20 vs 23.71), which is attributed to the highly structured and mask-friendly layout of LOGO-L, providing strong mask priors. Importantly, our model is mask-free and shows consistently better structural accuracy across all LOGO datasets (e.g., RMSE ↓1.6, SSIM ↑0.002, LPIPS ↓0.011 over Inpaint Anything on LOGO-H). These results confirm that DiffDecompose not only generalizes across our six diverse datasets but also outperforms existing methods on public benchmarks, without requiring additional spatial priors.